OpenStudy (anonymous):

Bryan’s monthly electric bill is determined by adding a flat administration fee to the product of the number of kilowatt hours of electricity used and the cost per kilowatt hour. When he uses 1,100 kilowatt hours of electricity, his bill is $113. When he uses 1,500 kilowatt hours of electricity, his bill is $153. What is the monthly administration fee?

OpenStudy (anonymous):

@Haseeb96

OpenStudy (haseeb96):

as u know @justsmile531 Bryan's monthly electric bill is determined by adding a flat administration fee to the product of the number of kilowatt hours of electricity used and the cost per kilowatt hour. okay let me take administration fee as ''x'' here bryan's monthly electric bill = x + ( number of kilowatt hours of electricity multiply with cost per kilowatt hour ) okay

OpenStudy (anonymous):

Ok , Now what

OpenStudy (mathmath333):

let the monthly administration fee be \(x\) and let the cost per kilowatt hour be \(y\) there fore \(\large\tt \begin{align} \color{black}{1100\times y=113-x }\end{align}\) and \(\large\tt \begin{align} \color{black}{1500\times y=153-x }\end{align}\) \(\large\tt \begin{align} \color{black}{\dfrac{1100\times y}{1500\times y}=\dfrac{113-x}{153-x}\\~\\ \dfrac{11}{15}=\dfrac{113-x}{153-x}\\~\\ \dfrac{15}{11}=\dfrac{153-x}{113-x}\\~\\ \dfrac{15-11}{11}=\dfrac{153-x-(113-x)}{113-x}\\~\\ \dfrac{4}{11}=\dfrac{40}{113-x}\\~\\ \dfrac{1}{11}=\dfrac{10}{113-x}\\~\\ 113-x=110\\~\\ \huge x=3 }\end{align}\)

OpenStudy (haseeb96):

Sorry i could not solve your question because it is out of my knowledge

OpenStudy (anonymous):

Thank you guys so much :)